# How Does A Text Preprocessing Pipeline Affect Ontology Syntactic Matching?

### v3

@@
Abstract
The generic text preprocessing pipeline, comprising Tokenisation, Normalisation, Stop Words Removal, and Stemming/Lemmatisation, has been implemented in many ontology matching (OM) systems. However, the lack of standardisation in text preprocessing creates diversity in mapping results. In this paper, we investigate the effect of the text preprocessing pipeline on OM tasks at syntactic levels. Our experiments on 8 Ontology Alignment Evaluation Initiative (OAEI) track repositories with 49 distinct alignments indicate: (1) Tokenisation and Normalisation are currently more effective than Stop Words Removal and Stemming/Lemmatisation; and (2) The selection of Lemmatisation and Stemming is task-specific. We recommend standalone Lemmatisation or Stemming with post-hoc corrections. We find that (3) Porter Stemmer and Snowball Stemmer perform better than Lancaster Stemmer; and that (4) Part-of-Speech (POS) Tagging does not help Lemmatisation. To repair less effective Stop Words Removal and Stemming/Lemmatisation used in OM tasks, we propose a novel context-based pipeline repair approach that significantly improves matching correctness and overall matching performance. We also discuss the use of text preprocessing pipeline in the new era of large language models (LLMs).
Index Terms:
@@ Abstract
I Introduction
Ontology matching (OM), also known as ontology alignment, is essential to enable interoperability between heterogeneous ontologies. An OM process usually takes two ontologies as input, discovers mappings between entities, and produces a set of correspondences [1]. A classical OM system usually contains syntactic, lexical, and semantic matching. This multilayer architecture has been implemented in several successful OM systems, such as LogMap [2][3], AgreementMakerLight (AML) [4][5], and FCA-Map [6][7].
@@ I Introduction
Syntactic matching captures “anchor mappings”, providing the foundation for the latter lexical and semantic matching. There are many strategies to extract syntactic information from an ontology entity, including the older approach of Bag-of-Words (e.g. TF-IDF [8]), popular word embedding models (e.g. Word2Vec [9]), and state-of-the-art language models (e.g. BERT [10]). Despite the diversity of the models used, they all apply text preprocessing for cleaning the text data before fitting it into the model. Fig. 1 shows an example of using a classical text preprocessing pipeline to process the ontology entity “cmt:reviewerBiddingStartedBy”. The text preprocessing pipeline consists of a set of steps to segment, reconstruct, analyse, and process the information in the text, namely Tokenisation, Normalisation, Stop Words Removal, and Stemming/Lemmatisation [11]. Tokenisation is the process of breaking the text into the smallest units (i.e. tokens). We use whitespace to split the tokens in the example. Normalisation is the process of transforming these different tokens into a single canonical form. Stop Words Removal is the process of removing filler words that usually carry little meaning and can be omitted in most cases. Stemming/Lemmatisation is used to deal with the grammatical variation of words, applying rules to find the simplest common form of the word. This helps to capture the key information from the text and therefore improves efficiency.
Figure 1: Text preprocessing pipeline used in syntactic matching.
@@
While a number of OM systems use the text preprocessing pipeline for syntactic OM, few studies explain why a specific method is selected for a certain OM task. Our study tackles the challenge in two ways. Firstly, we conduct a comprehensive experimental analysis of the text preprocessing pipeline in syntactic OM across a wide range of domains, aiming to explain the behaviour of the text preprocessing pipeline in OM tasks at syntactic levels. In each phase, a text preprocessing method is evaluated for its correctness and completeness. Secondly, we propose a novel context-based pipeline repair approach for syntactic OM. The method offers a customised way to fine-tune the text preprocessing pipeline for each domain-specific OM task and shows promising results for repairing false mappings. Thirdly, we discuss the non-deterministic nature of LLMs for text preprocessing and highlight the importance of retaining the classical text preprocessing pipeline. Specifically, this paper makes the following contributions:
@@
  • (Section IV) We categorise the text preprocessing pipeline used in syntactic OM into two phases. We find a significant improvement using Phase 1 text preprocessing methods. In contrast, Phase 2 text preprocessing methods are currently less effective. We compare the performance of (1) Stemming and Lemmatisation, (2) different stemming methods (Porter, Snowball, and Lancaster), and (3) Lemmatisation and Lemmatisation + Part-of-Speech (POS) Tagging. We find that inappropriate stop words removal, over-stemming, and over-lemmatisation are common on 8 Ontology Alignment Evaluation Initiative (OAEI) [12] track repositories with 49 distinct alignments.
@@
  • (Section V) We propose a simple and intuitive context-based pipeline repair method. This method is evaluated on the same OM tasks we analysed, showing promising results to improve the correctness of syntactic OM when inserted in the pipeline repair before Phase 2 text preprocessing methods.
  • (Section VI) We discuss the use of text preprocessing pipeline in the era of LLMs. We find LLMs cannot perform syntactic matching without text preprocessing, but they can be used as an alternative to repair the false mappings.
@@
The remainder of the paper is organised as follows: Section II and Section III review the related work and the preliminaries of OM task formulation. Section IV analyses the text preprocessing pipeline used in OM. Section V proposes the context-based pipeline repair approach and experimentally validates its performance. Section VI demonstrates the pros and cons of using LLMs for text processing in OM. Section VII and Section VIII summarise the findings and limitations of this work. Section IX concludes the paper.
II Related Work
@@
Syntactic matching considers only the meaning of the entity’s name or label, ignoring its lexical and structural context in an ontology [13]. Correct syntactic matches are often implicit and usually require extra human observation and domain knowledge. Text preprocessing is introduced to automate this process. The use of text preprocessing in syntactic matching can be traced back to the early stages of OM systems, having been initially developed and widely used as a basic component to generate linguistic-based mappings. AgreementMaker [14] has a normaliser to unify the textual information of the entities. SAMBO [15] uses the Porter Stemmer for each word to improve the similarity measure for terms with different prefixes and suffixes. In RiMOM [16], the context information of each entity is viewed as a document. The text in each document is preprocessed with tokenisation, stop words removal, and stemming. Recently, machine learning (ML) models have emerged for modern OM systems. While text preprocessing remains useful, its role is more focused on normalising the text that becomes the input to the model. For example, BERTMap [17] uses BERT’s inherent WordPiece tokeniser [18] to build the subword of each entity. A more recent approach DeepOnto [19] extends the normalisation to axioms using EL embedding models [20]. The ML extension of LogMap [21] reuses the seed mappings of the traditional system, where each entity is split into its component English word and the mapping is based on their similarity.
However, to the best of our knowledge, most of the literature implements a preprocessing method without explaining why a specific method is chosen, and no studies have been conducted to evaluate the effect of text preprocessing on syntactic OM.
III Preliminaries: OM task formulation
@@
Given a source ontology $O_{s}$ and a target ontology $O_{t}$, OM establishes mappings between pairs of entities drawn from each of two ontologies. A correspondence (i.e. one instance of mappings) is defined as a 4-tuple $(e_{1},e_{2},r,c)$, where $e_{1}\in O_{s}$ and $e_{2}\in O_{t}$. $r$ is the relationship between two matched entities $e_{1}$ and $e_{2}$, and $c\in[0,1]$ is the confidence for each correspondence. The relationship $r$ in OM tasks can be equivalence ($\equiv$), subsumption ($\subseteq$), disjointness ($\perp$), or other more complex relationships. In this paper, we focus only on the equivalence relationship ($\equiv$) and evaluate the effect of text preprocessing on syntactic matching to produce the “anchor mappings” on which to base any subsequent lexical and semantic matching. We address only equivalence because subsumption and disjointness are typically dealt with in a later semantic and structural matching phase of OM. An alignment (A) is a set of candidate correspondences generated by tools, while a Reference (R) is a set of gold standard correspondences verified by domain experts (i.e. the ground truth alignment).
IV Text Preprocessing Pipeline: Analysis
@@
In this section, we discuss the role of the text preprocessing pipeline in ontology syntactic matching and report our experiments studying the effect of different text preprocessing options over the OAEI repositories.
IV-A Method
@@
Fig. 2 shows the method used to analyse the text preprocessing pipeline. Alignment (A) is generated via the text preprocessing pipeline. Firstly, for both $O_{s}$ and $O_{t}$, we retrieve the entities from the named classes (i.e. owl:Class) and named properties (i.e. object properties owl:ObjectProperty and data type properties owl:DatatypeProperty). For those ontologies where the names of the concepts are not textual (e.g. a numerical identifier), instead, we retrieve the meaningful text from entity labels (i.e. rdfs:label). Then, we apply the text preprocessing pipeline method $f(.)$ on each entity $e_{1}\in O_{s}$ and $e_{2}\in O_{t}$. If $f(e_{1})=f(e_{2})$, we store the correspondence in the corresponding alignment file. Finally, we compare the generated Alignment (A) with Reference (R) to evaluate the performance of the text preprocessing pipeline on OM tasks.
@@
Figure 2: Method used to analyse the text preprocessing pipeline. B-Base Entity without Text Preprocessing, T-Tokenisation, N-Normalisation, R-Stop Words Removal, S/L-Stemming/Lemmatisation.
@@
    Selected OAEI track repositories: The Ontology Matching Evaluation Toolkit (MELT) [22] is a powerful framework for OM evaluation. We retrieve 17 OAEI normal track repositories stored in the MELT public repository. The repositories are categorised as schema matching or instance matching. 12 of the 17 track repositories are schema-matching and applicable for evaluating OM. Only 8 of these track repositories are selected because the other 4 have noisy data or miss required files. Specifically, the knowledgegraph repository contains both schema and instance matching. The complex repository is primarily focused on detecting complex correspondences, while the multifarm repository is an extension of the conference repository to multilingualism. The laboratory repository lacks reference files at the time of writing. Table I shows the details of the selected track repositories. We consider only the equivalence mappings contained in the reference files. Each track repository may contain more than one alignment corresponding to different pairs of ontologies. For example, the largebio track repository has 6 reference files, pairing the whole and small versions of FMA [23] and NCI [24], FMA and SNOMED [25], and SNOMED and NCI, respectively. The number of references for each track repository is given in the table. There are 49 distinct alignments evaluated in this study.
    Name Domain Number of References anatomy Human and Mouse Anatomy 1 biodiv Biodiversity and Ecology 9 commonkg Common Knowledge Graphs 3 conference Conference 24 food Food Nutritional Composition 1 largebio Biomedical 6 mse Materials Science & Engineering 3 phenotype Disease and Phenotype 2
    TABLE I: Selected OAEI track repositories.
    Compound words are frequently used in ontology naming conventions. For example, compound words “art gallery” can be formatted as the Pascal case “ArtGallery” used in YAGO [26] or the Snake case “art_gallery” used in Wikidata [27]. Fig. 3 shows the proportion of compound words in the selected OAEI track repositories. For comparisons between entities with compound words, one approach is to use their tokens, where “XY” is equivalent to “YX” because they share the same concepts X and Y. However, our comparison considers both concepts and their order, so that “XY” is not equivalent to “YX”. Although concurrent “XY” and “YX” are rare in the matching process, we assume that it could happen and so we take account of the order of concepts in compound word names to ensure a fair comparison in our experiments.
    Figure 3: The proportion of compound words.
@@
    Selected subtasks of pipeline methods: There are a variety of subtasks that can be used in a general text preprocessing pipeline, but not all of them are applicable to OM tasks. We select the following subtasks in each text preprocessing method: (1) Tokenisation: includes word tokenisation only. We do not use sentence tokenisation (also known as segmentation) because text retrieved from the entity’s name or label is commonly short. (2) Normalisation: includes lowercasing, HTML tags removal, separator formatting, and punctuation removal. Other subtasks that may potentially change the word semantics, such as removal of special characters and numbers, are excluded. (3) Stop Words Removal: includes the most common English stop words defined in the Natural Language Toolkit (NLTK) [28]. (4) Stemming/Lemmatisation: Stemming methods include Porter Stemmer, Snowball Stemmer, and Lancaster Stemmer. Lemmatisation uses the NLTK Lemmatiser [28] based on WordNet [29], and the word categorisation uses POS Tagging.
@@
    Selected OM evaluation measures: In information retrieval, there are four primitive measures: true positive (TP), false positive (FP), false negative (FN), and true negative (TN). In the context of OM, evaluation compares an alignment (A) returned by the OM system with a gold standard reference (R). Fig. 4 illustrates that the four primitive measures in OM can be interpreted as $TP=A\cap R$, $FP=A-R$, $FN=R-A$, and $TN=(C\times C^{\prime})-(A\cup R)$, where $C\times C^{\prime}$ refers to all possible correspondences $\in\{O_{s},O_{t}\}$.
    Figure 4: OM evaluation measures [30].
@@
    Accuracy (Acc), Specificity (Spec), Precision (Prec), Recall (Rec), and $F_{\beta}$ Score ($F_{\beta}$) are the most common evaluation measures based on TP, FP, FN, and TN. In the context of OM, since $C\times C^{\prime}$ is extremely large (the Cartesian product of $e_{1}\in O_{s}$ and $e_{2}\in O_{t}$), TN is often much larger than TP, FP, and FN. This means that Accuracy (Acc) and Specificity (Spec) are close to 1, and they have no statistically significant difference across different alignments. We note that Precision (Prec) and Recall (Rec) contribute equally to $F_{\beta}$. Therefore, we choose Precision (Prec), Recall (Rec), and F1 Score ($\beta=1$) in this study. They are defined as:
    $\footnotesize Prec=\frac{|A\cap R|}{|A|}\quad Rec=\frac{|A\cap R|}{|R|}\quad F_{1}=\frac{2}{Prec^{-1}+Rec^{-1}}$ (1)
IV-B Results

### v4

# How Does A Text Preprocessing Pipeline Affect Syntactic Ontology Matching?

@@
Abstract.
The generic text preprocessing pipeline, comprising Tokenisation, Normalisation, Stop Words Removal, and Stemming/Lemmatisation, has been implemented in many systems for syntactic ontology matching (OM). However, the lack of standardisation in text preprocessing creates diversity in mapping results. In this paper, we investigate the effect of the text preprocessing pipeline on syntactic OM in 8 Ontology Alignment Evaluation Initiative (OAEI) tracks with 49 distinct alignments. We find that Phase 1 text preprocessing (Tokenisation and Normalisation) is currently more effective than Phase 2 text preprocessing (Stop Words Removal and Stemming/Lemmatisation). To repair the less effective Phase 2 text preprocessing caused by unwanted false mappings, we propose a novel context-based pipeline repair approach that employs an ad hoc check to find common words that cause false mappings. These words are stored in a reserved word set and applied in text preprocessing. The experimental results show that our approach improves the matching correctness and the overall matching performance. We also discuss the integration of the classical text preprocessing pipeline with modern large language models (LLMs). We recommend that LLMs inject the text preprocessing pipeline via function calling to avoid the tendency towards unstable true mappings produced by prompt-based LLM approaches, and use LLMs to repair false mappings generated by the text preprocessing pipeline.
@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: 29th International Conference on Extending Database Technology; March 24–27, 2026; Tampere, Finlandisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Ontology engineering
1. Introduction
@@
Ontology matching (OM) is essential for enabling interoperability between heterogeneous ontologies. An OM process usually takes two ontologies as input, discovers mappings between entities, and produces a set of correspondences (Euzenat and Shvaiko, 2013). A classical OM system usually contains syntactic matching, lexical matching, and semantic matching. This multilayer architecture has been implemented in several successful OM systems, such as LogMap (Jiménez-Ruiz and Cuenca Grau, 2011; Jiménez-Ruiz et al., 2011), AgreementMakerLight (AML) (Faria et al., 2013, 2014), and FCA-Map (Zhao et al., 2018; Li et al., 2021). While syntactic matching captures “anchor mappings” and provides the foundation for the latter lexical and semantic matching, there are many strategies for extracting syntactic information from an ontology entity. These include older approaches Bag-of-Words (e.g. TF-IDF (Sammut and Webb, 2010)), popular word embedding models (e.g. Word2Vec (Mikolov et al., 2013)), and state-of-the-art language models (e.g. BERT (Devlin et al., 2019)). Despite the diversity of the models used, they all apply the text preprocessing pipeline to clean the text data before fitting them into the model. The pipeline consists of a set of steps to segment, reconstruct, analyse, and process the information in the text, namely Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L) (Manning et al., 2008). Table 1 illustrates examples of using the text preprocessing pipeline to find mappings in syntactic OM. T is the process of breaking the text into the smallest units (i.e. tokens). N is the process of transforming these different tokens into a single canonical form. R is the process of removing filler words that usually carry little meaning and can be omitted in most cases. S/L is used to deal with the grammatical variation of words, applying rules to find the simplest common form of the word.
Pipeline Ontology Entity Name Output T cmt ProgramCommittee Program Committee sigkdd Program_Committee N cmt ProgramCommittee program committee conference Program_committee R conference has_a_review review ekaw hasReview S/L conference reviewes review edas isReviewing
Table 1. Examples of using the text preprocessing pipeline in syntactic OM: Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L). In the examples, we use whitespace to split the tokens, normalise all the words into lowercase, remove the common stop words “has” and “a”, and apply S/L to find the simplest word “review” from “reviewes” and “reviewing”.
@@
While a number of OM systems use the text preprocessing pipeline for syntactic OM, few studies explain why a specific text preprocessing method is selected. Our study tackles this challenge. Firstly, we conduct a comprehensive experimental analysis of the text preprocessing pipeline in syntactic OM across a wide range of domains, aiming to explain the behaviour of the text preprocessing pipeline in syntactic OM. In each step, text preprocessing methods are evaluated for correctness and completeness. Secondly, we propose a novel context-based pipeline repair approach for syntactic OM. The method offers a customised way to fine-tune the text preprocessing pipeline for each domain-specific OM and shows promising results for repairing less effective text preprocessing methods. Thirdly, we discuss the non-deterministic nature of modern large language models (LLMs) for text preprocessing and highlight the importance of integrating the classical text preprocessing pipeline with LLMs. Specifically, this paper makes the following contributions:
@@
  • (Section 4) We compare the performance of the text preprocessing pipeline in 8 Ontology Alignment Evaluation Initiative (OAEI) (OAEI Community, [n. d.]) datasets with 49 distinct alignments. While there is a significant improvement using Tokenisation (T) and Normalisation (N) in syntactic OM, Stop Words Removal (R) and Stemming/Lemmatisation (S/L) are currently less effective. We find inappropriate stop words removal, over-stemming, and over-lemmatisation. We also compare the performance of several text preprocessing methods: (1) Stemming and Lemmatisation, (2) different stemming methods (Porter, Snowball, and Lancaster), and (3) Lemmatisation with and without Part-of-Speech (POS) Tagging.
@@
  • (Section 5) We propose a simple and intuitive context-based pipeline repair method using a reserved word set to capture certain words that may lead to false mappings. This method is evaluated on the same OAEI datasets we analysed, showing promising results in improving the matching correctness and the overall matching performance of syntactic OM.
  • (Section 6) We discuss the integration of the classical text preprocessing pipeline with modern LLMs. We recommend using text preprocessing via function calling to overcome the non-deterministic of purely prompt-based approaches in finding true mappings, while LLMs show incredible potential to repair false mappings generated by the text preprocessing pipeline.
@@
The remainder of the paper is organised as follows. Section 2 and 3 review the related work and the preliminaries of OM task formulation. Section 4 analyses the text preprocessing pipeline used in syntactic OM. Section 5 proposes the context-based pipeline repair approach and experimentally validates its performance. Section 6 demonstrates the integration of the classical text preprocessing pipeline with modern LLMs in syntactic OM. Section 7 and 8 discuss the findings and limitations. Section 9 concludes the paper.
2. Related Work
@@
Syntactic OM considers only the meaning of the entity’s name or label, ignoring its lexical and structural context in an ontology. Correct syntactic matches are often implicit and usually require extra human observation and domain knowledge (Euzenat and Shvaiko, 2013). Text preprocessing is introduced to automate this process. The use of text preprocessing in syntactic OM can be traced back to the early stages of OM systems, which were initially developed and widely used as a basic component to generate linguistic-based mappings. AgreementMaker (Cruz et al., 2009) has a normaliser to unify the textual information of the entities. SAMBO (Lambrix and Tan, 2006) uses the Porter Stemmer for each word to improve the similarity measure for terms with different prefixes and suffixes. In RiMOM (Li et al., 2009), the context information of each entity is viewed as a document. The text in each document is preprocessed with tokenisation, stop words removal, and stemming. Recently, machine learning (ML) models have emerged for modern OM systems. While text preprocessing remains useful, its role is more focused on normalising the text that becomes the input to the model. For example, BERTMap (He et al., 2022) uses BERT’s inherent WordPiece tokeniser to build the subword of each entity. DeepOnto (He et al., 2024) extends normalisation to axioms using EL embedding models (Kulmanov et al., 2019). The ML extension of LogMap (Chen et al., 2021) reuses the seed mappings of the traditional system, where each entity is split into its component word and the mapping is based on their similarity. However, to the best of our knowledge, most of the literature implements a preprocessing method without explaining why a specific method is chosen, and no studies have been conducted to evaluate the effect of the entire text preprocessing pipeline on syntactic OM.
3. Preliminaries: OM Task Formulation
@@
Given a source ontology ($O_{s}$) and a target ontology ($O_{t}$), OM establishes mappings between pairs of entities drawn from each of the two ontologies. A correspondence (i.e. one instance of mappings) is defined as a 4-tuple $(e_{1},e_{2},r,c)$, where $e_{1}\in O_{s}$ and $e_{2}\in O_{t}$. $r$ is the relationship between two matched entities $e_{1}$ and $e_{2}$, and $c\in[0,1]$ is the confidence for each correspondence. The relationship $r$ in OM tasks can be equivalence ($\equiv$), subsumption ($\subseteq$), or other more complex relationships. In this study, we focus only on the equivalence relationship ($\equiv$) and evaluate the effect of text preprocessing on syntactic matching to produce the “anchor mappings” on which to base any subsequent lexical and semantic matching. We address only equivalence because subsumption is typically dealt with in a later semantic and structural matching phase of OM. An alignment (A) is a set of candidate correspondences generated by tools, while a reference (R) is a set of gold standard correspondences verified by domain experts (i.e. the ground truth alignment).
4. Analysis of Text Preprocessing Pipeline
@@
In this section, we discuss the role of the text preprocessing pipeline in syntactic OM and report our experiments that study the effect of different text preprocessing options over the OAEI datasets.
4.1. Experiment Setup
@@
Figure 1 shows the experiment setup to analyse the text preprocessing pipeline. Alignment (A) is generated via the text preprocessing pipeline. For both $O_{s}$ and $O_{t}$, we retrieve the entities from classes (owl:Class) and properties (both owl:ObjectProperty and owl:DatatypeProperty). For those ontologies where the names of the concepts are not textual (e.g. a code or numerical identifier), we retrieve the meaningful text from entity labels (rdfs:label or other annotation properties) instead. Then, we apply the text preprocessing pipeline method $f(\cdot)$ on each entity $e_{1}\in O_{s}$ and $e_{2}\in O_{t}$. If $f(e_{1})=f(e_{2})$, we store the correspondence in the corresponding alignment file. We compare the generated alignment (A) with the reference (R) to evaluate the performance of the text preprocessing pipeline on syntactic OM.
@@
Figure 1. Experiment setup to analyse the text preprocessing pipeline: Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L).
@@
    Selected OAEI tracks: We select 8 OAEI tracks stored in the Ontology Matching Evaluation Toolkit (MELT) (Hertling et al., 2019) public repository (retrieved January 1, 2025). Table 2 shows the details of the selected tracks. Each track may contain more than one alignment corresponding to different pairs of ontologies. For example, the Largebio Track has multiple reference files, pairing FMA (Rosse and Mejino, 2003) and NCI (Golbeck et al., 2003), FMA and SNOMED (Donnelly, 2006), and SNOMED and NCI, respectively. The number of alignments for each track is given in the table. There are 49 distinct alignments evaluated in this study.
    Table 2. Selected OAEI tracks and the number of alignments.
    Name Domain Number of Alignments Anatomy Human and Mouse Anatomy 1 Biodiv Biodiversity and Ecology 9 CommonKG Common Knowledge Graphs 3 Conference Conference 24 Food Food Nutritional Composition 1 Largebio Biomedical 6 MSE Materials Science & Engineering 3 Phenotype Disease and Phenotype 2
    Compound words are frequently used in ontology naming conventions. For example, compound words such as “art gallery” can be formatted as the Pascal case “ArtGallery” used in YAGO (Suchanek et al., 2007) or the Snake case “art_gallery” used in Wikidata (Vrandečić and Krötzsch, 2014). Figure 2 shows the proportion of compound words in the selected OAEI tracks. For comparisons between entities with compound words, one approach is to use their tokens, where “XY” is equivalent to “YX” because they share the same concepts X and Y. However, our comparison considers both concepts and their order, so that “XY” is not equivalent to “YX”. Although concurrent “XY” and “YX” are rare in the matching process, we assume that it could happen and so we take account of the order of concepts in compound word names to ensure a fair comparison in our experiments.
    Figure 2. Proportion of compound words.
@@
    Selected pipeline methods: There are a variety of methods that can be used in a general text preprocessing pipeline, but not all of them are applicable to syntactic OM. We select the following methods in each text preprocessing step: (a) Tokenisation (T): We use full word tokenisation to separate the compound words. Experiments on several subword tokenisation methods (e.g. Byte Pair Encoding (Gage, 1994; Sennrich et al., 2016), WordPiece (Schuster and Nakajima, 2012; Devlin et al., 2019), and Unigram (Kudo, 2018; Kudo and Richardson, 2018)) show the results after de-tokenisation (i.e. converting the tokens to a sentence) are generally the same as full word tokenisation. We do not use sentence tokenisation (a.k.a. segmentation) because text retrieved from the entity’s name or label is commonly short. (b) Normalisation (N): includes lowercasing, HTML tags removal, separator formatting, and punctuation removal. Other methods that may potentially change the word semantics are excluded (e.g. removal of special characters and numbers). (c) Stop Words Removal (R): includes the most common English stop words defined in the Natural Language Toolkit (NLTK) (Bird et al., 2008). (d) Stemming/Lemmatisation (S/L): Stemming methods include Porter, Snowball, and Lancaster. Lemmatisation uses the NLTK Lemmatiser based on WordNet (Miller, 1995), and the word categorisation uses POS Tagging.
@@
    Selected measures: There are four primitive measures in information retrieval: true positive (TP), false positive (FP), false negative (FN), and true negative (TN). In the context of OM, the evaluation compares an alignment (A) returned by the OM system with a gold standard reference (R) validated by domain experts. Figure 3 illustrates that the four primitive measures in OM can be interpreted as $TP=A\cap R$, $FP=A-R$, $FN=R-A$, and $TN=(C\times C^{\prime})-(A\cup R)$, where $C\times C^{\prime}$ refers to all possible correspondences $\in\{O_{s},O_{t}\}$. Accuracy (Acc), Specificity (Spec), Precision (Prec), Recall (Rec), and $F_{\beta}$ Score ($F_{\beta}$) are the most common evaluation measures based on TP, FP, FN, and TN. In the context of OM, since $C\times C^{\prime}$ is extremely large (the Cartesian product of $e_{1}\in O_{s}$ and $e_{2}\in O_{t}$), TN is often much larger than TP, FP, and FN. This means that Accuracy (Acc) and Specificity (Spec) are close to 1, and they have no statistically significant difference across different alignments. We note that Precision (Prec) and Recall (Rec) contribute equally to $F_{\beta}$. Therefore, we choose Precision (Prec), Recall (Rec), and F1 Score ($\beta=1$) in this study. They are defined as follows.
    Figure 3. OM Evaluation (Euzenat, 2007).
    $Precision(Prec)=\frac{|A\cap R|}{|A|}$
    $Recall(Rec)=\frac{|A\cap R|}{|R|}$
    $F_{1}\ Score=\frac{2}{Prec^{-1}+Rec^{-1}}$
4.2. Results
4.2.1. Comparison of Text Preprocessing Pipeline
Figure 4 shows the comparison of the text preprocessing pipeline. These include Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L). The methods are always applied sequentially in the pipeline. For T and N, most of the data points above the equivalent line in precision, recall, and F1 score indicate that they can help syntactic OM. For R and S/L, some data points are above the equivalent line in recall, but most data points are below the equivalent line in precision and F1 score indicate that they do not play a prime role. Three horizontal lines inside each violin plot show three quartiles: Q1, median, and Q3. The extension of the curves beyond 0% and 100% is an artefact of the seaborn (Waskom, 2021) package. In the violin plots: (a) Precision: The median increases with T and N but decreases with R and S/L. After T, the shape of the distribution is unchanged by R and S/L. (b) Recall: After T, the median increases slightly with each of N, R, and S/L. The shape of the distribution does not change after N. (c) F1 Score: The median increases with T and N but then decreases with R and S/L. The shape of the distribution does not change after N.
Figure 4. Comparison of text preprocessing pipeline: Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L). The symbol $\varnothing$ represents the baseline without text preprocessing.
4.2.2. Stemming vs. Lemmatisation
Figure 5 shows the comparison of Stemming (S) and Lemmatisation (L) after Tokenisation (T) and Normalisation (N) have been applied. Most data points above the L = S line in precision and F1 score indicate that L is better than S in syntactic OM (assuming that post hoc correction is excluded). In the violin plots: (a) Precision: The median after L is greater than that achieved by S. The shape of the distribution is slightly different. (b) Recall: The median and the shape of the distribution are identical after S and after L. (c) F1 Score: The median after L is greater than for S. The shape of the distribution is identical.
Figure 5. Stemming (S) vs. Lemmatisation (L).
4.2.3. Different Stemming Methods
Figure 6 shows the comparison of different stemming methods: Porter Stemmer (SP), Snowball Stemmer (SS), and Lancaster Stemmer (SL). For SP and SS, all data points located on the equivalent line indicate that there is no difference between SP and SS. For SP/SS and SL, most data points above the equivalent line in precision and F1 score indicate that SP/SS is more effective than SL. In the violin plots: (a) Precision: The median number of SP and SS is greater than that of SL, and there is no difference between SP and SS. The shape of the distribution is identical. (b) Recall: The median number and the shape of the distribution are identical in SP, SS, and SL. (c) F1 Score: The median number of SP and SS is greater than that of SL, and there is no difference between SP and SS. The shape of the distribution is identical in SP, SS, and SL.
Figure 6. Comparison of different stemmers: Porter Stemmer (SP), Snowball Stemmer (SS), and Lancaster Stemmer (SL).
4.2.4. Lemmatisation with/without POS Tagging
Figure 7 shows the comparison of Lemmatisation (L) and Lemmatisation + POS Tagging (LT). All data points located on the equivalent line indicate that using POS Tagging in Lemmatisation does not help syntactic OM. In the violin plots, for each of (a) Precision, (b) Recall, and (c) F1 Score, the median number and the shape of the distribution are identical for L and LT.
Figure 7. Lemmatisation (L) without/with POS Tagging (LT).
4.3. Discussion
For syntactic OM, the text preprocessing pipeline can be categorised into two phases. Phase 1 text preprocessing methods contain Normalisation and Tokenisation, whereas Phase 2 preprocessing methods consist of Stop Words Removal and Stemming/Lemmatisation. Phase 1 text preprocessing methods do not change word semantics, but change syntactic features such as formatting and typography. In contrast, Phase 2 text preprocessing methods change words to better capture semantic similarity, such as removing prefixes and suffixes or substituting common etymological roots.